# Neural Solver Selection for Combinatorial Optimization

### v1

@@ Abstract
Machine learning has increasingly been employed to solve NP-hard combinatorial optimization problems, resulting in the emergence of neural solvers that demonstrate remarkable performance, even with minimal domain-specific knowledge. To date, the community has created numerous open-source neural solvers with distinct motivations and inductive biases. While considerable efforts are devoted to designing powerful single solvers, our findings reveal that existing solvers typically demonstrate complementary performance across different problem instances. This suggests that significant improvements could be achieved through effective coordination of neural solvers at the instance level. In this work, we propose the first general framework to coordinate the neural solvers, which involves feature extraction, selection model, and selection strategy, aiming to allocate each instance to the most suitable solvers. To instantiate, we collect several typical neural solvers with state-of-the-art performance as alternatives, and explore various methods for each component of the framework. We evaluated our framework on two extensively studied combinatorial optimization problems, Traveling Salesman Problem (TSP) and Capacitated Vehicle Routing Problem (CVRP). Experimental results show that the proposed framework can effectively distribute instances and the resulting composite solver can achieve significantly better performance (e.g., reduce the optimality gap by 0.88% on TSPLIB and 0.71% on CVRPLIB) than the best individual neural solver with little extra time cost. Our code is available at https://github.com/lamda-bbo/neural-solver-selection.
@@ 1 Introduction
To enhance the capabilities of neural solvers, a variety of methods have been proposed, with intensive effort on the design of frameworks, network architectures, and training procedures. For example, to improve the performance across different distributions, Jiang et al. (2022) proposed adaptively joint training over varied distributions, and Bi et al. (2022) leveraged knowledge distillation to integrate the models trained on different distributions. For generalization on large-scale instances, Fu et al. (2021) implemented a divide-and-conquer strategy, Luo et al. (2023) proposed a heavy-decoder structure to better capture the relationship among nodes, while Gao et al. (2024) utilized the local transferability and introduced an additional local policy model. Diffusion models (Sun & Yang, 2023) have also been adapted to generate the distribution of optimal solutions, demonstrating impressive results. More works include bisimulation quotienting (Drakulic et al., 2023), latent space search (Chalumeau et al., 2023), local reconstruction (Cheng et al., 2023; Ye et al., 2024; Zheng et al., 2024) and so on.
@@
Figure 1: (a), (b): Observation from the comparison of prevailing neural solvers at instance level. Details of the settings are provided in Section 4.1. (c): Our proposed coordination framework.
As various neural solvers are emerging in the community, the state-of-the-art records for the overall performance on benchmark problems are frequently refreshed. However, the detailed comparison of these neural solvers on each instance has been rarely discussed. Here, we empirically examined the performance of several prevailing neural solvers on each instance, as illustrated in Figure 1(a) and 1(b). As expected from the no-free-lunch theorem (Wolpert & Macready, 1997), we find that
  • As shown in Figure 1(a), there exists no single neural solver that can dominate all the other neural solvers on every instance, and different neural solvers win on different instances, demonstrating their complementary performance at instance level.
  • As shown in Figure 1(b), the modification of instance distribution can almost reverse the domination relationship of neural solvers, which further verifies that different neural solvers are good at instances with specific characteristics due to their intrinsic inductive biases.
@@
Motivated by the observations above, we, for the first time, propose a general framework to coordinate end-to-end neural solvers for COPs at instance level by selecting suitable individual solvers for each instance, as illustrated in Figure 1(c). Specifically, our proposed framework consists of three key components, which are summarized as follows:
  • Feature extraction: For each problem instance, extract features that can be effectively used to identify their characteristics.
  • Selection model: Based on the features of instances, train a selection model that can be utilized to identify suitable solvers for each instance.
  • Selection strategy: Due to the intricate structures of COPs, using only the most suitable individual solver predicted by the selection model may fail. Therefore, it is important to design robust selection strategies based on the confidence of the selection model.
To verify the effectiveness of our proposed framework, we collect several prevailing open-source neural solvers and their released models with competitive performance in the community to construct the pool of individual solvers, and provide several implementations for each component of the framework. For feature extraction, we utilize the graph attention network (Veličković et al., 2018; Kool et al., 2019) to encode COP instances, and further propose a refined encoder with pooling to leverage the hierarchical structures of COPs. For selection model, we train it from the perspective of classification and ranking, respectively. We also implement several selection strategies, including top-$k$ selection, rejection-based selection, and so on. Detailed descriptions are provided in Section 3. We conduct experiments on two widely studied COPs: Traveling Salesman Problem (TSP) and Capacitated Vehicle Routing Problem (CVRP). Experimental results exhibit that our framework can generally select suitable individual solvers for each instance to achieve significantly better performance with limited extra time consumption. Compared to the best individual solver, our framework reduces the optimality gap by 0.82% on synthetic TSP, 2.01% on synthetic CVRP, 0.88% on TSPLIB (Reinelt, 1991), and 0.71% on CVRPLIB Set-X (Uchoa et al., 2017). As this is the first preliminary attempt on neural solver selection for COPs, we also analyze the influence of various implementations of components, and provide some discussion on future improvements.
@@ 2.1 End-to-end neural solvers for COPs
Traditional approaches for COPs have achieved impressive results, but they often rely on problem-specific heuristics and domain knowledge by experts (Helsgaun, 2000; 2017). Instead, recent efforts focus on utilizing end-to-end learning methods. A prominent fashion is autoregression, which employs graph neural networks in an encoder-decoder framework and progressively extends a partial solution until a complete solution is constructed (Vinyals et al., 2015; Bello et al., 2017; Kool et al., 2019). However, these methods tend to exhibit poor generalization performance across distributions and scales (Joshi et al., 2022). To address the generalization issue, considerable efforts have been dedicated within the community. For example, Zhou et al. (2023) took various distributions and scales as different learning tasks and adopted meta-learning over them to obtain a generalizable model. Bi et al. (2022) leveraged knowledge distillation, where models trained on different distributions are utilized as teacher models for one generalizable student model. Liu et al. (2024) used the idea of prompt learning to realize zero-shot adaptation of the pertained model by selecting the most matched prompt for instances. More efforts in autoregressive methods include instance-conditioned adaptation (Zhou et al., 2024a), adversarial training (Wang et al., 2024) and nested local views (Fang et al., 2024), to name a few.
Another popular kind of end-to-end learning methods is non-regressive, which predicts or generates the distributions of potential solutions. Typically,  Joshi et al. (2019); Ye et al. (2023) employed graph neural networks to predict the probability of components appearing in an optimal solution, represented with the form of heatmap. Diffusion models (Sun & Yang, 2023; Sanokowski et al., 2024) have also been adapted to generate the distribution of optimal solutions, demonstrating better expressiveness than classical push-forward generative models (Salmona et al., 2022).
@@ 2.2 Solving COPs with multiple neural solvers
Recent studies have made preliminary attempts to integrate multiple neural solvers to enhance overall performance on COPs. For example, Jiang et al. (2023) adopted ensemble learning, where multiple neural solvers with identical architecture are trained on different instance distributions through Bootstrap sampling to ensure diversity. During inference, the outputs of all the solvers are gathered by average at each action step. Grinsztajn et al. (2023) proposed a population-based training method Poppy, where multiple decoders with a shared encoder are trained simultaneously as a population of solvers, with a reward targeting at maximizing the overall performance of the population. When solving a problem instance, each solver generates solutions independently, and the best solution is selected as the final result. However, these works suffer from heavy computation cost as multiple solvers have to be run for each instance. Even they propose to share a common encoder for each solver, experimental results still demonstrate undesired inference time (Grinsztajn et al., 2023). On the other hand, different solvers share the same neural architecture, which may limit the diversity and thus the final performance.
Consider that the burgeoning community has proposed many methods from various perspectives, resulting in diverse end-to-end neural solvers with different inductive biases. Properly coordinating these neural solvers can potentially bring a significant improvement on overall performance. Motivated by the observation in Figure 1(a) and 1(b), we propose to select suitable ones from a pool of diverse individual solvers for each instance. Note that similar idea has been utilized in the area of algorithm selection (Kerschke et al., 2019) and model selection (Zhang et al., 2023), but has never been explored in the area of neural combinatorial optimization. By solver selection at instance level, any type of (existing or newly constructed) neural solver can be utilized, and only the selected individual solvers need to be run in inference, thereby maintaining high efficiency.
@@ 3 The proposed framework
This section will introduce the proposed framework of coordinating neural solvers for COPs. In general, our target is learning to select the suitable solvers for each problem instance. To address this, our proposed framework comprises three key components:
  • Feature extraction: To select the most suitable neural solvers for each instance, it is essential to extract the instance features, which is challenging as the COPs are usually intricate. In this work, we first utilize the graph attention encoder (Kool et al., 2019) to encode COP instances, and further propose a refined graph encoder with pooling, which can leverage the hierarchical structures of COPs to obtain better features.
@@
  • Selection strategies: Due to the complexity of COPs, it may be risky to rely solely on the selection model to consistently identify the most suitable solver. To address this, we propose compromise strategies that allow to allocate multiple solvers (if necessary) to a single instance based on the confidence levels of the selection model, pursuing better performance with limited extra cost.
In the following subsections, we will elaborate the three key components in our framework.
@@ 3.1 Feature extraction
Graph attention encoder.
We take the CVRP as an example to describe the computation of the graph encoder. The raw features $\bm{x}\in\mathbb{R}^{N\times 3}$ of a CVRP instance are a set of nodes $\{(x_{i},y_{i},m_{i})|i\in[N]\}$, where $(x_{i},y_{i})$ are the node coordinates, $m_{i}$ is the node demand, $N$ is the number of nodes, and $[N]$ denotes the set $\{1,2,\ldots,N\}$. First, a linear layer is employed on every node for initial node embeddings, i.e., $H^{0}=\bm{x}W$, where $W\in\mathbb{R}^{3\times d}$ are the weights and $d$ denotes the embedding dimension. Given initial embeddings, multiple graph attention layers (Veličković et al., 2018; Kool et al., 2019) are applied to iteratively update the node embeddings as $H^{l}=\text{AttentionLayer}^{l}(H^{l-1})$, where $l\in[L]$ and $L$ is the number of layers. Since the graphs of TSP and CVRP are both fully connected, the graph attention layer covers every pair of nodes and self-connections, which becomes similar to the self-attention mechanism (Vaswani et al., 2017). Details of the attention layer are provided in Appendix A.1. Finally, the node embeddings output by the last layer are averaged to form the instance representation, as most COP encoders (Khalil et al., 2017; Kool et al., 2019) did.
Hierarchical graph encoder.
Averaging the final node embeddings may result in sub-optimal instance representations that are too flat to effectively capture the hierarchical structures inherent in COPs (Goh et al., 2024). Inspired by (Lee et al., 2019), we design a hierarchical graph attention encoder to address this limitation, which successively downsamples the graph of an instance using graph pooling, and aggregates features from each downsampling level to construct a comprehensive graph representation, as illustrated in Figure 2.
@@
The hierarchical graph encoder contains $L$ blocks. In each block, several graph attention layers are first applied, and then the graph pooling layer selects representative nodes to form a coarsened graph that preserves important features. Consider the $l$-th block, where the number of selected nodes is denoted as $N_{l}$. To quantify the representativeness of each node for graph pooling, an additional graph attention layer is introduced to compute representative scores. Specifically, this graph attention layer computes score embeddings $H^{l}_{\text{score}}$ based on the current node embeddings $H^{l}$, which encode rich information about the graph structure and node features. These score embeddings $H^{l}_{\text{score}}$ are then mapped to scalar representative scores via linear layer. The complete process is shown as follows:
@@
where $Z^{l}\in\mathbb{R}^{N^{l-1}\times 1}$ are the representative scores of the $N^{l-1}$ nodes preserved in the $(l-1)$-th block, $\sigma$ is a non-linear function (here we use the $\tanh$ function), and $W^{l}_{\text{score}}\in\mathbb{R}^{d\times 1}$ are the parameters of the linear layer. Subsequently, we sort the nodes according to their representative scores and select top-$N^{l}$ nodes ($N^{l}<N^{l-1}<N$) to preserve. To make this pooling layer trainable via back-propagation (LeCun et al., 2002), we further combine the representative score together with the embeddings of their corresponding nodes as follows: $\tilde{H}^{l}=H^{l}+Z^{l}\bm{1}$, where $\bm{1}\in\mathbb{R}^{1\times d}$ is a vector with all the elements being $1$. Intuitively, this operation can move the embeddings of high scored nodes away from the embeddings of low scored nodes.
Figure 2 shows the complete encoder, where $L$ blocks are stacked, and each block is formed by several graph attention layers followed by a pooling layer. By successively applying $L$ encoder blocks, the number of preserved nodes gradually decreases according to $N^{l}=\alpha\cdot N^{l-1}$ ($\alpha$ is set to 0.8 in our experiments). This process constructs a hierarchy of the original graph and its coarsened versions, enabling the encoder to capture multi-level structural information effectively. Within each block, we apply a readout layer that aggregates the embeddings after the graph attention layers by mean pooling and max pooling (Lee et al., 2019), i.e.,

### v2

@@ Abstract
Machine learning has increasingly been employed to solve NP-hard combinatorial optimization problems, resulting in the emergence of neural solvers that demonstrate remarkable performance, even with minimal domain-specific knowledge. To date, the community has created numerous open-source neural solvers with distinct motivations and inductive biases. While considerable efforts are devoted to designing powerful single solvers, our findings reveal that existing solvers typically demonstrate complementary performance across different problem instances. This suggests that significant improvements could be achieved through effective coordination of neural solvers at the instance level. In this work, we propose the first general framework to coordinate the neural solvers, which involves feature extraction, selection model, and selection strategy, aiming to allocate each instance to the most suitable solvers. To instantiate, we collect several typical neural solvers with state-of-the-art performance as alternatives, and explore various methods for each component of the framework. We evaluated our framework on two typical problems, Traveling Salesman Problem (TSP) and Capacitated Vehicle Routing Problem (CVRP). Experimental results show that our framework can effectively distribute instances and the resulting composite solver can achieve significantly better performance (e.g., reduce the optimality gap by 0.88% on TSPLIB and 0.71% on CVRPLIB) than the best individual neural solver with little extra time cost. Our code is available at https://github.com/lamda-bbo/neural-solver-selection.
Machine Learning, ICML
@@ 1 Introduction
To enhance the capabilities of neural solvers, a variety of methods have been proposed, with intensive effort on the design of frameworks, network architectures, and training procedures. For example, to improve the performance across different distributions, (Jiang et al., 2022) proposed adaptively joint training over varied distributions, and (Bi et al., 2022) leveraged knowledge distillation to integrate the models trained on different distributions. For generalization on large-scale instances, (Fu et al., 2021) implemented a divide-and-conquer strategy, (Luo et al., 2023) proposed a heavy-decoder structure to better capture the relationship among nodes, while (Gao et al., 2024) utilized the local transferability and introduced an additional local policy model. Diffusion models (Sun & Yang, 2023) have also been adapted to generate the distribution of optimal solutions, demonstrating impressive results. More works include bisimulation quotienting (Drakulic et al., 2023), latent space search (Chalumeau et al., 2023), local reconstruction (Cheng et al., 2023; Ye et al., 2024; Zheng et al., 2024) and so on.
@@
Figure 1: (a), (b): Observation from the comparison of prevailing neural solvers at instance level. Details of the settings are provided in Section 4.1. (c): Our proposed selection framework.
As various neural solvers are emerging in the community, the state-of-the-art records for the overall performance on benchmark problems are frequently refreshed. However, the detailed comparison of these neural solvers on each instance has been rarely discussed. Here, we empirically examined the performance of several prevailing neural solvers on each instance, as illustrated in Figure 1(a) and 1(b). As expected from the no-free-lunch theorem (Wolpert & Macready, 1997), we find that:
$\bullet$ As shown in Figure 1(a), there exists no single neural solver that can dominate all the others on every instance, and different neural solvers win on different instances, demonstrating their complementary performance at instance level.
$\bullet$ As shown in Figure 1(b), the modification of instance distribution can almost reverse the domination relationship of neural solvers, which further verifies that different neural solvers are good at instances with specific characteristics due to their intrinsic inductive biases.
@@
Motivated by the observations above, we, for the first time, propose a general framework to coordinate end-to-end neural solvers for COPs at the instance level by selecting suitable individual solvers for each instance, as illustrated in Figure 1(c). Specifically, our proposed framework consists of three key components, which are summarized as follows:
$\bullet$ Feature extraction: For each problem instance, extract features for effectively identifying their characteristics.
$\bullet$ Selection model: Based on the features of instances, train a selection model that can be utilized to identify suitable solvers for each instance.
$\bullet$ Selection strategy: Due to the intricate structures of COPs, using only the most suitable individual solver predicted by the selection model may fail. Therefore, it is important to design robust selection strategies based on the confidence of the selection model.
To verify the effectiveness of our proposed framework, we collect several prevailing open-source neural solvers and their released models with competitive performance in the community to construct the pool of individual solvers, and provide several implementations for each component of the framework. For feature extraction, we utilize the graph attention network (Veličković et al., 2018; Kool et al., 2019) to encode COP instances, and further propose a refined encoder with pooling to leverage the hierarchical structures of COPs. For selection model, we train it from the perspective of classification and ranking, respectively. We also implement several selection strategies, including top-$k$ selection, rejection-based selection, and so on. Detailed descriptions are provided in Section 3. We conduct experiments on two widely studied COPs: Traveling Salesman Problem (TSP) and Capacitated Vehicle Routing Problem (CVRP). Experimental results exhibit that our framework can generally select suitable individual solvers for each instance to achieve significantly better performance with limited extra time consumption. Compared to the best individual solver, our framework reduces the optimality gap by 0.82% on synthetic TSP, 2.00% on synthetic CVRP, 0.88% on TSPLIB (Reinelt, 1991), and 0.71% on CVRPLIB Set-X (Uchoa et al., 2017). As the first preliminary attempt on neural solver selection for COPs, we also analyze the influence of various implementations of components, and discuss on future directions.
@@ 2.1 End-to-end neural solvers for COPs
Traditional approaches for COPs have achieved impressive results, but they often rely on problem-specific heuristics and domain knowledge by experts (Helsgaun, 2000, 2017). Instead, recent efforts focus on utilizing end-to-end learning methods. A prominent fashion is autoregression, which employs graph neural networks in an encoder-decoder framework and progressively extends a partial solution until a complete solution is constructed (Vinyals et al., 2015; Bello et al., 2017; Kool et al., 2019). However, these methods tend to exhibit poor generalization performance across distributions and scales (Joshi et al., 2022). To address the generalization issue, considerable efforts have been dedicated from various perspectives, including meta-learning (Zhou et al., 2023), knowledge distillation (Bi et al., 2022), prompt learning (Liu et al., 2024), instance-conditioned adaptation (Zhou et al., 2024a), adversarial training (Wang et al., 2024) and nested local views (Fang et al., 2024).
Another popular kind of end-to-end learning methods is non-regressive, which predicts or generates the distributions of potential solutions. Typically,  (Joshi et al., 2019; Ye et al., 2023) employed graph neural networks to predict the probability of components appearing in an optimal solution, represented with the form of heatmap. Diffusion models (Sun & Yang, 2023; Sanokowski et al., 2024) have also been adapted to generate the distribution of optimal solutions, demonstrating better expressiveness than classical push-forward generative models (Salmona et al., 2022).
@@ 2.2 Solving COPs with multiple neural solvers
Recent studies have made preliminary attempts to integrate multiple neural solvers to enhance overall performance on COPs. For example, (Jiang et al., 2023) adopted ensemble learning, where multiple neural solvers with identical architecture are trained on different instance distributions through Bootstrap sampling. During inference, the outputs of all the solvers are averaged at each action step. (Grinsztajn et al., 2023) proposed a population-based training method Poppy, where multiple decoders with a shared encoder are trained simultaneously as a population of solvers, with a reward targeting at maximizing the overall performance of the population. When solving a problem instance, each solver generates solutions independently, and the best solution is selected as the final result. However, these works suffer from heavy cost as multiple solvers have to be run for each instance. Even they propose to share a common encoder for each solver, experiments still demonstrate undesired inference time (Grinsztajn et al., 2023). On the other hand, different solvers share the same neural architecture, which may limit the diversity and thus the final performance.
Consider that the burgeoning community has proposed many methods from various perspectives, resulting in diverse neural solvers with different inductive biases. Properly coordinating them can potentially bring significant improvement on overall performance. Motivated by the observation in Figure 1(a) and 1(b), we propose to select suitable ones from a pool of diverse individual solvers for each instance. Note that similar idea has been utilized in the area of algorithm selection (Kerschke et al., 2019; He et al., 2025) and model selection (Zhang et al., 2023), but has never been explored in the area of neural combinatorial optimization. By solver selection at instance level, any existing or newly constructed neural solver can be utilized, and only the selected individual solvers need to be run in inference, thereby maintaining high efficiency.
@@ 3 The proposed framework
This section introduces our proposed framework of coordinating neural solvers for COPs. Our target is learning to select the suitable solvers for each problem instance. To address it, the framework comprises three key components:
Feature extraction: To select the most suitable neural solvers for each instance, it is essential to extract the instance features, which is challenging as the COPs are usually intricate. In this work, we first utilize the graph attention encoder (Kool et al., 2019) to encode COP instances, and further propose a refined graph encoder with pooling, which can leverage the hierarchical structures of COPs.
@@
Selection strategies: Due to the complexity of COPs, it may be risky to rely solely on the selection model to identify the best solver. To address this, we propose compromise strategies to allocate multiple solvers (if necessary) to a single instance based on the confidence levels of the selection model, pursuing better performance with limited extra cost.
@@ 3.1 Feature extraction
Graph attention encoder. We take the CVRP as an example to describe the computation of the graph encoder. The raw features $\bm{x}\in\mathbb{R}^{N\times 3}$ of a CVRP instance are a set of nodes $\{(x_{i},y_{i},m_{i})|i\in[N]\}$, where $(x_{i},y_{i})$ are the node coordinates, $m_{i}$ is the node demand, $N$ is the number of nodes, and $[N]$ denotes the set $\{1,2,\ldots,N\}$. First, a linear layer is employed on every node for initial node embeddings, i.e., $H^{0}=\bm{x}W$, where $W\in\mathbb{R}^{3\times d}$ are the weights and $d$ denotes the embedding dimension. Given initial embeddings, multiple graph attention layers (Veličković et al., 2018; Kool et al., 2019) are applied to iteratively update the node embeddings as $H^{l}=\text{AttentionLayer}^{l}(H^{l-1})$, where $l\in[L]$ and $L$ is the number of layers. Since the graphs of TSP and CVRP are both fully connected, the graph attention layer covers every pair of nodes and self-connections, which becomes similar to the self-attention mechanism (Vaswani et al., 2017). Details of the attention layer are in Appendix A.1. Finally, the node embeddings output by the last layer are averaged to form the instance representation, like most COP encoders (Khalil et al., 2017; Kool et al., 2019).
Hierarchical graph encoder. Averaging the final node embeddings may result in sub-optimal instance representations that are too flat to effectively capture the hierarchical structures inherent in COPs (Goh et al., 2024).
@@
Inspired by (Lee et al., 2019), we design a hierarchical graph attention encoder to address this limitation, which successively downsamples the graph of an instance using graph pooling, and aggregates features from each downsampling level to construct a comprehensive graph representation, as illustrated in Figure 2. The hierarchical graph encoder contains $L$ blocks. In each block, several graph attention layers are applied, and then the graph pooling layer selects representative nodes to form a coarsened graph that preserves important features. Consider the $l$-th block, where the number of selected nodes is denoted as $N_{l}$. To quantify the representativeness of each node for graph pooling, an additional graph attention layer is introduced to compute representative scores. Specifically, the graph attention layer computes score embeddings $H^{l}_{\text{score}}$ based on the current node embeddings $H^{l}$, which encode rich information about the graph structure and node features. These score embeddings $H^{l}_{\text{score}}$ are then mapped to scalar representative scores via linear layer. The complete process is shown as follows:
@@
where $Z^{l}\in\mathbb{R}^{N^{l-1}\times 1}$ are the representative scores of the $N^{l-1}$ nodes preserved in the $(l-1)$-th block, $\sigma$ is a non-linear function (here we use the $\tanh$ function), and $W^{l}_{\text{score}}\in\mathbb{R}^{d\times 1}$ are the parameters of the linear layer. Subsequently, we sort the nodes according to their representative scores and select top-$N^{l}$ nodes ($N^{l}<N^{l-1}<N$) to preserve. To make this pooling layer trainable via back-propagation (LeCun et al., 2002), we further combine the representative score together with the embeddings of their corresponding nodes as follows: $\tilde{H}^{l}=H^{l}+Z^{l}\bm{1}$, where $\bm{1}\in\mathbb{R}^{1\times d}$ is a vector with all the elements being $1$. Intuitively, this operation can separate the embeddings of high scored nodes from the embeddings of low scored nodes.
Figure 2 shows the complete encoder, where $L$ blocks are stacked, and each block is formed by several graph attention layers followed by a pooling layer. By successively applying $L$ encoder blocks, the number of preserved nodes gradually decreases as $N^{l}=\alpha\cdot N^{l-1}$ ($\alpha$ is set to 0.8 in our experiments). This process constructs a hierarchy of the original graph and its coarsened versions, enabling the encoder to capture multi-level structural information effectively. Within each block, we apply a readout layer that aggregates the embeddings after the graph attention layers by mean pooling and max pooling (Lee et al., 2019), i.e.,